Let x and y be real numbers. Prove that, if 5x+y>11, then x>2 or y>1.


1
Expert's answer
2021-11-23T09:24:54-0500

for x = y:

"6y>11\implies y>11/6 \implies y>1"


for x > y:

"x=y+k,k>0"

"6x-k>11\implies x>(11+k)/6"

"6y+5k>11"

if "y\le 1" then "k>1\implies x>2"


for x < y:

"x=y-k,k>0"

"6x+k>11"

"6y-5k>11\implies y>11/6+5k/6>1"


so, if 5x+y>11, then x>2 or y>1